The cooling system in a truck contains of 4 gallon of coolant that is 30% antifreeze. How much must be withdrawn and replaced with 100% antifreeze to bring the coolant in the system to 50% antifreeze?
Answer by sofiyac(983)

let's say we need x gallons of 100% antifreeze so
4(30)-x(40)+x(100)=4(50)
120-40x+100x=200
60x=80 divide each side by 60
x=8/6=4/3 of a gallon or 1.33333 gallons